OpenStudy (chillhill):

Mick invested 1500 francs at an annual rate of interest of 5.25% compounded annually. What should the interest rate be if Michele's initial investment to double in value in 10 years?

OpenStudy (anonymous):

well the simple form of the compound interest formula is x(1+r)^t so since we want it to be doubled we can write it as 3000=1500(1+r)^10

OpenStudy (anonymous):

if we divide both sides by 1500 we get 2 = (1+r)^10 if we take the 10th root of both sides and subtract 1, we get \[r = \pm \sqrt[10]{2} -1\]

OpenStudy (anonymous):

since \[-\sqrt[10]{2} -1 = -2.07177..\] and \[\sqrt[10]{2}-1 = 0.071773462\] 0.07177 is the correct choice.